OpenStudy (anonymous):

how do I isolate the y variable in this equation. 2x+y=1

jimthompson5910 (jim_thompson5910):

Subtract 2x from both sides to isolate y

OpenStudy (anonymous):

the answer is -1

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

2x+y=1 2x+y - 2x=1 - 2x y = -2x + 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

think I get it

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

so x-y=5 to isolate the y I would add x to both sides right

jimthompson5910 (jim_thompson5910):

To solve for y, subtract x from both sides x - y = 5 -y = 5 - x Keep going to completely isolate y

OpenStudy (anonymous):

why was that so hard. I get it! Thank you!!!

jimthompson5910 (jim_thompson5910):

yw, tell me what you get for the final answer

OpenStudy (anonymous):

y=5+x

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

close, but it's really y = -5+x or y = x - 5

OpenStudy (anonymous):

I always forget to put that part

OpenStudy (anonymous):

So in my first equation, I need to continue to solve for y right